# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

### v1

@@ Abstract
Key Words and Phrases: symmetric equation, existence of solutions, equivariant Leray-Schauder degree, nonlinear analysis, wave equation.
@@ 1 Introduction
One of the paradigmatic problems used to validate novel techniques in nonlinear analysis is classical forced pendulum equation $\ddot{x}=\bm{f}(x)$ (cf. [7]-[11]). More generally, dynamical systems described by second-order differential equations strike a delicate balance between simplicity and a capacity for capturing rich and intricate nonlinear behavior, making them ideal candidates for the testing of analytical methods. On the other hand, autonomous differential equations admit natural $SO(2)$ symmetries related to time shifting that make them particularly well-suited for a framework based on the equivariant degree theory. Given a finite group $\Gamma$ and a fixed number $p>0$, our primary objective is to demonstrate how the Leray-Schauder equivariant degree can be used to establish the existence of non-stationary solutions to the $\Gamma$-symmetric network of wave equations
@@
    the Jacobian of $f$ at the origin $Df(0):V\rightarrow V$ is trivial, i.e. for all $x\in V$ one has
    $Df(0)x=0.$
@@
In the context of an arrangement of $N$ identical pendula, subjected to nonlinear forcing represented by $f$ and with coupling relations described by the matrix $A$, the vectors
$\begin{cases}x(t)=(x_{1}(t),x_{2}(t),\ldots,x_{N}(t)),\\ \dot{x}(t)=(\dot{x}_{1}(t),\dot{x}_{2}(t),\ldots,\dot{x}_{N}(t));\\ \ddot{x}(t)=(\ddot{x}_{1}(t),\ddot{x}_{2}(t),\ldots,\ddot{x}_{N}(t)),\end{cases}$
contain the angles of the pendula from the vertical position, their velocities and their accelerations at any time $t>0$, respectively. The effectiveness of the equivariant Leray-Schauder degree in detecting solutions to a system of a similar form to (1) for several explicit choices of the symmetry group $\Gamma$ has already been demonstrated by Garcia-Azpetia et al. in the article [5]. Our contribution to this problem is a practical existence result that applies for any arbitrary choice of $\Gamma$ and our technique is broadly applicable to any model exhibiting symmetries of the product group
$G:=O(2)\times\Gamma\times{\mathbb{Z}}_{2}.$
1.1 Setting for the Leray-Schauder Equivariant Degree
With the substitutions $\beta:=p/2\pi$ and $u(t):=x(\beta t)$, the equations (1) become
$\begin{cases}\ddot{u}(t)=\beta^{2}f(u(t))+\beta^{2}Au(t),\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+2\pi),\;\dot{u}(t)=\dot{u}(t+2\pi).\end{cases}$ (2)
Notice that any $2\pi$-periodic solution $u(t):\mathbb{R}\rightarrow V$ for the system (2) corresponds to the $p$-periodic solution $x(t)=u(t/\beta)$ for our original system (1). We denote by $\mathscr{E}:=C_{2\pi}(\mathbb{R},V)$ the Banach space of all continuous, $2\pi$-periodic, $V$-valued functions with the standard norm
$\|\varphi\|_{\infty}:=\max_{t\in\mathbb{R}}|\varphi(t)|,\quad\varphi\in C_{2\pi}(\mathbb{R},V),$
and by $\mathscr{H}:=C^{2}_{2\pi}(\mathbb{R},V)$ the subspace of twice continuously differentiable functions equipped with the norm
$\|u\|_{\mathscr{H}}:=\max\{\|u\|_{\infty},\|\dot{u}\|_{\infty},\|\ddot{u}\|_{\infty}\},\quad u\in\mathscr{H}.$
Following [5], we also consider the differential operator
the Nemystskii operator
$N:\mathscr{E}\rightarrow\mathscr{E},\quad(Nu)(t):=f(u(t)),$
and the Banach embedding
$j:\mathscr{H}\rightarrow\mathscr{E},\quad j(u)(t):=u(t).$
Since $\mathscr{L}:\mathscr{H}\rightarrow\mathscr{E}$ is a linear isomorphism, $N:\mathscr{E}\rightarrow\mathscr{E}$ is continuous and $j:\mathscr{H}\rightarrow\mathscr{E}$ is compact, the operator $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ given by
$\mathscr{F}(u)=0,$ (4)
in the sense that a function $u\in\mathscr{H}$ is a solution to (2) if and only if it satisfies (4). In what follows, (2) will be called the operator equation associated with the system (2). Notice that the space $\mathscr{H}$ is an isometric Banach $G$-representation with respect to the $G$-action $G\times\mathscr{H}\rightarrow\mathscr{H}$
$(e^{i\theta},\gamma,\pm 1)u(t):=\pm\gamma u(t+\theta),\quad(\kappa,\gamma,\pm 1)u(t):=\pm\gamma u(-t),\quad u\in\mathscr{H},\;t\geq 0.$
where $(e^{i\theta},\gamma,\pm 1),(\kappa,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$. Gracia-Azpetia et al. demonstrate the applicability of the $G$-equivariant Leray-Schauder degree to (4) in [5] by proving that the assumptions ($A_{1}$)($A_{4}$) imply that nonlinear operator (3) is a completely continuous $G$-equivariant field, differentiable at the origin $0\in\mathscr{H}$ with
$\mathscr{A}u=\operatorname{Id}-\mathscr{L}^{-1}(\beta^{2}Au-u),\quad\mathscr{A}:=D_{u}\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}.$
For proof of the following Lemma, the reader is referred to [2, 3].
@@ 1.1 Setting for the Leray-Schauder Equivariant Degree
Under the conditions ($A_{1}$)-($A_{4}$), the operator $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ is $\Omega:=B_{R}(0)\setminus B_{\epsilon}(0)$-admissible for some sufficiently small choice of $\epsilon>0$ and a sufficiently large choice of $R>0$. Moreover, one has
$G\text{\rm-deg}(\mathscr{F},\Omega)=(G)-G\text{\rm-deg}(\mathscr{A},B(\mathscr{H})),$
where $(G)\in A(G)$ is the unit element of the Burnside Ring.
Lemma (1.1) provides a sufficient condition for the existence of a non-trivial solution to (2); namely, if one has
$G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\neq(G),$ (5)
then there exists a function $u\in\mathscr{H}\setminus\{0\}$ satisfying (2). A more practical reformulation of the condition (5): if for some orbit type $(H)\in\Phi_{0}(G)\setminus\{(G)\}$ one has
then, by the existence property of the $G$-equivariant Leray-Schauder degree (cf. Appendix A), there exists a map $u\in\mathbb{R}$ with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
Without specifying the group $\Gamma$, we assume that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=0}^{r}$ is made available. As a $\Gamma$-representation, the space $V$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
@@ 1.2 Towards a Computational Formula for $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
In order to simplify our analysis, we enforce an additional non-degeneracy assumption:
@@
Remark 1.2.
In [5], Garcia-Azpetia et al. avoid the assumption ($A_{0}$) with an appropriate fixed point reduction. For the sake of simplicity, we choose to sidestep the fixed point argument by assuming ($A_{0}$).
For each $s\in{\mathbb{N}}$, we denote by $\mathcal{W}_{s}\simeq\mathbb{C}$ the irreducible $O(2)$-representation equipped with the $s$-folding $O(2)$-action
$e^{i\theta}w:=e^{is\theta}w,\quad\kappa w:=\bar{w},\quad\theta\in SO(2),\;\kappa\in O(2),\;w\in\mathcal{W}_{s},$
and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation on which $O(2)$ acts trivially. Notice that the space $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
@@
In the standard way, the $G$-isotypic decomposition of $\mathscr{H}$ can now be expressed in terms of $G$-isotypic components $\mathscr{H}_{m,j}\subset\mathscr{H}$ modeled on the irreducible $G$-representations
@@
$\mathscr{H}=\bigoplus_{j=0}^{r}\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m,j}},\quad\mathscr{H}_{m,j}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V_{j}\}.$
With the $G$-isotypic decomposition of $\mathscr{H}$ at hand, we can begin to collect spectral data related to the $G$-equivariant linear operator $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$. For example, by Schur’s Lemma, one has
$\mathscr{A}(\mathscr{H}_{m,j})\subset\mathscr{H}_{m,j},\quad m\in{\mathbb{N}}\cup\{0\},\;j\in\{0,1,\ldots,r\}.$

### v2

@@ Abstract
Key Words and Phrases: symmetric equation, existence of solutions, equivariant Leray-Schauder degree, nonlinear analysis, periodic solutions.
@@ 1 Introduction
The classical forced pendulum equation $\ddot{x}=\bm{f}(x)$ serves as a paradigmatic problem for validating novel techniques in nonlinear analysis (cf. [20]-[24]). More generally, dynamical systems described by second-order differential equations strike a delicate balance between simplicity and a capacity for capturing rich and intricate nonlinear behavior, making them ideal candidates for the testing of analytical methods. On the other hand, autonomous differential equations with periodic boundary conditions always admit natural $SO(2)$ symmetries related to time shifting that make them particularly well-suited for a framework based on the equivariant degree theory. Second order autonomous differential equations not involving any first derivative terms, specifically, admit $O(2)$ symmetries, where the reflection generator acts by time reversal. When these equations are subjected to an odd forcing function and arranged in symmetric networks characterized by a finite group $\Gamma$, their full symmetry group becomes
$G:=O(2)\times\Gamma\times{\mathbb{Z}}_{2}.$
Our primary objective is to demonstrate how the $G$-equivariant Leray-Schauder degree can be used to establish the existence of non-stationary solutions to the following $\Gamma$-symmetric system of $p$-periodic, second order, autonomous equations:
@@
    $f(x)$ is $o(|x|)$ as $x$ approaches 00, i.e.
    $\lim_{x\rightarrow 0}\frac{f(x)}{|x|}=0.$
@@
Conditions ($A_{1}$)($A_{2}$) ensure that the equations admit the symmetries of the product group $O(2)\times\Gamma\times{\mathbb{Z}}_{2}$, while conditions ($A_{3}$) and ($A_{4}$) imply the well-behavedness of our non-linearity at solutions to (1) near zero and infinity, respectively. Specifically, ($A_{2}$) allows the equivariant degree to distinguish between stationary and non-stationary functions, ($A_{3}$) guarantees that the nonlinearity does not interact with the spectrum of $A$ and the Nagumo condition ($A_{4}$) is essential for establishing a-priori bounds on the magnitude of solutions to (1). System (1) is a natural model for the dynamics of a configuration of $N$ identical pendula, subjected to nonlinear forcing $f$ and with coupling relations specified by the matrix $A$ and the group $\Gamma$. In this context, the vector $x(t):=(x_{1}(t),x_{2}(t),\ldots,x_{N}(t))$ represents the angles of the pendula from the vertical position, $\dot{x}(t):=(\dot{x}_{1}(t),\dot{x}_{2}(t),\ldots,\dot{x}_{N}(t))$ the angular velocities and $\ddot{x}(t):=(\ddot{x}_{1}(t),\ddot{x}_{2}(t),\ldots,\ddot{x}_{N}(t))$ the angular accelerations at any time $t>0$. Methods based on the Leray-Schauder degree have been used to solve a wide variety of second order differential equations [2, 9, 14, 17, 18, 22, 24]. However, the inability of the Leray-Schauder degree to distinguish between periodic solutions that differ by a fixed time shift and differentiate constant solutions from non-constant solutions have hindered its suitibility for autonomous differential equations. The effectiveness of the $G$-equivariant Leray-Schauder degree in detecting non-constant solutions to the system (1) for several explicit choices of the symmetry group $\Gamma$ has already been demonstrated by Garcia-Azpetia et al. in the article [13]. More generally, the $G$-equivariant Leray-Schauder degree has been applied to problems admitting appropriate spatial-temporal symmetries [3, 4, 11, 12, 15, 19]. Compared with other techniques to solve systems similar to (1), the equivariant degree approach requires relatively minimal regularity assumptions on the nonlinearity. For example, existence results were obtained in [1] by topological index argument and in [10] using Morse theory framework, under the assumption that the right-hand-side interacts with only finitely many eigenvalues of the differential operator. A prerequisite for the application of any $G$-equivariant degree-theory based argument to (1) is the reformulation of our system as a fixed point equation in an appropriate functional $G$-space with a nonlinear operator in the form of a $G$-equivariant compact perturbation of identity. In [13], Garcia-Azpetia et. al prove that the problem of finding $p$-periodic solutions to (1) is equivalent to the problem of finding zeros of the operator
$\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H},\quad\mathscr{F}(u):=u-\mathscr{L}^{-1}(\beta^{2}f(u)+\beta^{2}Au-u),$
where $\beta:=\frac{p}{2\pi}$, $\mathscr{H}:=H^{2}_{2\pi}([0,2\pi];V)$ and $\mathscr{L}:\mathscr{H}\rightarrow\mathscr{H}$ is the differential operator $\mathscr{L}u:=\ddot{u}-u$. Under the conditions ($A_{1}$)($A_{4}$) and provided that the linearization $D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism, Garcia-Azpetia et. al additionally demonstrate, that:
  1. $\rm(i)$
    $\mathscr{F}$ is a $G$-equivariant completely continuous field with respect to the natural isometric action of $G$ on $\mathscr{H}$;
  2. $\rm(ii)$
    all of the non-trivial solutions to (1) can be confined to an annulus of the form $\Omega:=B_{R}(\mathscr{H})\setminus B_{\varepsilon}(\mathscr{H})\subset\mathscr{H}$ for some sufficiently small choice of $\varepsilon>0$ and sufficiently large choice of $R>0$;
  3. $\rm(iii)$
    $(\mathscr{F},\Omega)$ constitutes an admissible $G$-pair, such that the existence of non-trivial solutions to (1) is equivalent to non-triviality of the degree invariant
    $G\text{\rm-deg}(\mathscr{F},\Omega)=(G)-G\text{\rm-deg}(D\mathscr{F}(0),B_{1}(\mathscr{H})),$
[…]
Together with a direct application of the existence property of the $G$-equivariant Leray-Schauder degree (cf. Appendix A), the above considerations suggest the following sufficient condition for the existence of a non-trivial solution to (1):
@@ 3. $\rm(iii)$
If for some orbit type $(H)\in\Phi_{0}(G)\setminus\{(G)\}$ one has
$\operatorname{coeff}^{H}\left(G\text{\rm-deg}(D\mathscr{F}(0),B(\mathscr{H}))\right)\neq 0,$
(cf. Appendix A for definition of the coefficient operator $\operatorname{coeff}^{H}:A(G)\rightarrow{\mathbb{Z}}$) then there exists a function $u\in\mathscr{H}\setminus\{0\}$ satisfying (4) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
In this way, the problem of finding $p$-periodic solutions to (1) has been reformulated as the problem of computing the coefficients of non-unit orbit types (i.e. orbit types $(H)\in\Phi_{0}(G)\setminus\{(G)\}$) in the Burnside Ring element $G\text{\rm-deg}(D\mathscr{F}(0),B(\mathscr{H}))$. Moreover, since the $G$-equivariant Leray-Schauder degree provides a full equivariant classification to the solution set of $\mathscr{F}(u)=0,\;u\in\mathscr{H}$, Lemma 1.1 represents a framework for identifying the exact spatio-temporal symmetries of any predicted solution. In order to take advantage of the full power of the $G$-equivariant Leray-Schauder degree for this purpose, we must first identify the irreducible representations of $G$ and describe the $G$-isotypic decomposition of $\mathscr{H}$. Without specifying the group $\Gamma$, we assume that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=0}^{r}$ is made available. As a $\Gamma$-representation, the space $V$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
@@
For the sake of simplicity, we choose to sidestep a fixed point argument made by Garcia-Azpetia et al. in [13] by enforcing the following non-degeneracy assumption:
@@
On the other hand, for each $m\in{\mathbb{N}}$, we denote by $\mathcal{W}_{m}\simeq\mathbb{C}$ the irreducible $O(2)$-representation equipped with the $m$-folding $O(2)$-action
$e^{i\theta}w:=e^{im\theta}w,\quad\kappa w:=\bar{w},\quad\theta\in SO(2),\;\kappa\in O(2),\;w\in\mathcal{W}_{m},$
and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation on which $O(2)$ acts trivially such that the space $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
@@
The $G$-isotypic decomposition of $\mathscr{H}$ can now be expressed in terms of $G$-isotypic components $\mathscr{H}_{m,j}\subset\mathscr{H}$ modeled on the irreducible $G$-representations
@@
$\mathscr{H}=\bigoplus_{j=0}^{r}\overline{\bigoplus\limits_{m=0}^{\infty}\mathscr{H}_{m,j}},\quad\mathscr{H}_{m,j}:=\{u\in\mathscr{H}:u(t)=\cos(mt)a+\sin(mt)b,\;a,b\in V_{j}\}.$ (3)
We are now in a position to present our main existence results:
Theorem 1.2.
Take $m>0$ and let $(H)$ be a maximal element in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$. Under the assumptions ($A_{0}$)($A_{4}$), if the matrix $A:V\rightarrow V$ admits an odd number of eigenvalues $\mu_{j}\in\sigma(A)$ satisfying
$\mu_{j}<-\left(\frac{m}{\beta}\right)^{2}\text{ and }2\nmid\dim\mathcal{V}_{m,j}^{H},$
then there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (1) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
Remark 1.3.
In the field of equivariant bifurcation theory, the Ize Conjecture (IC) refers to the proposition that every irreducible representation of a compact Lie group admits at least one subgroup with an odd dimensional fixed point space. Although the IC was shown to be false by Lauterbach et al. in [25] via counterexample, the conjecture holds in many cases. For example, in Section 7.1, we use G.A.P. to numerically verify that the irreducible representations of the group $G=O(2)\times\Gamma\times{\mathbb{Z}}_{2}$, with the assignment $\Gamma=D_{8}$, satisfy the IC.
The remainder of this paper is organized as follows: In Section 2, we prepare the problem (1) for application of the $G$-equivariant Leray-Schauder degree. This involves a functional reformulation in an appropriate Sobolev space, allowing us to derive a practical formula for a degree invariant, in terms of the Burnside Ring product of a finite number of basic degrees defined on the irreducible $G$-representations, whose non-triviality is equivalent to the existence of non-trivial solutions to (1). Using this framework, we analyze the maximal elements in the isotropy lattices $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$, called orbit types of maximal kind in $\Phi_{0}(G)$. The significance of these orbit types is two-fold: first, the coefficients corresponding to orbit types of maximal kind in the degree invariant are computationally accessible and second, the non-triviality of these coefficients guarantees the existence of non-stationary solutions to (1). In Section 4, we present a classification for the generators of the Burnside Ring $A(G)$ according to their lattice relations with respect to the $s$-folding homomorphism (13) and employ this classification, in Section 5, to derive a practical formula for characterizing the non-triviality of an orbit type of maximal kind in the product of a finite number of basic degrees. Specifically, we find that the contribution of each eigenvalue $\mu_{m,j}\in\sigma(D\mathscr{F}(0))$ to the appearance of an orbit type of maximal kind $(H)$ in the degree invariant $G\text{\rm-deg}(\mathscr{F},\Omega)$ depends on $(\rm i)$ the sign of $\mu_{m,j}$ and $(\rm ii)$ the parity of the dimension of the corresponding $H$-fixed-point space $\mathcal{V}_{m,j}^{H}$. Finally, in Section 7, we consider the implications of our main result for dihedral symmetry group $\Gamma=D_{N}$. For the special case of $D_{8}$ coupling symmetries, we use GAP to numerically verify $(\rm i)$ the pairwise disjointness of the maximal orbit type sets in each isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ and $(\rm ii)$ the odd-dimensionality of the fixed point sets $\mathcal{V}_{m,j}^{H}$ for every maximal element in $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$, allowing us to reformulate Theorem 1.2 in Section 7.1 as follows:
Proposition 1.4.
Let $\Gamma=D_{8}$ and take any fixed $m>0$. Under the assumptions ($A_{0}$)($A_{4}$), if the $j$-th eigenvalue of the matrix $A$ satisfies
$\mu_{j}<-\left(\frac{m}{\beta}\right)^{2},$
then, corresponding to every maximal element $(H)$ in the isotropy lattice $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$, there exists a non-stationary solution $u\in\mathscr{H}\setminus\{0\}$ to the system (1) with an isotropy subgroup $G_{u}\leq G$ satisfying $(G_{u})\geq(H)$.
For convenience, the Appendices include an explanation of notations used, and a brief introduction to the $G$-equivariant Leray-Schauder degree. Readers who are interested in a deeper exposition of these topics are referred to [6, 7].
2 Setting for the Leray-Schauder $G$-Equivariant Degree
Notice that, with the substitutions $\beta:=p/2\pi$ and $u(t):=x(\beta t)$, the system (1) become:
$\begin{cases}\ddot{u}(t)=\beta^{2}f(u(t))+\beta^{2}Au(t),\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+2\pi),\;\dot{u}(t)=\dot{u}(t+2\pi),\end{cases}$ (4)
such that any $2\pi$-periodic solution $u(t):\mathbb{R}\rightarrow V$ for the system (4) corresponds to the $p$-periodic solution $x(t)=u(t/\beta)$ for our original equation. Following [13], we denote by $\mathscr{H}$ the Sobolev space of $2\pi$-periodic, $V$-valued functions
:=$\mathscr{H}:=\{u:\mathbb{R}\rightarrow V:u(0)=u(2\pi),\;u|_{[0,2\pi]}\in H^{2}(S^{1};V)\},$
equipped with the standard inner product
[…]
and the isometric $G$-action
$(e^{i\theta},\gamma,\pm 1)u(t):=\pm\gamma u(t+\theta),\quad(\kappa,\gamma,\pm 1)u(t):=\pm\gamma u(-t),\quad u\in\mathscr{H},\;t\geq 0,$
where $(e^{i\theta},\gamma,\pm 1),(\kappa,\gamma,\pm 1)\in O(2)\times\Gamma\times{\mathbb{Z}}_{2}$. We also consider the differential operator
the Nemystskii operator
$N:L^{2}(S^{1};V)\rightarrow L^{2}(S^{1};V),\quad(Nu)(t):=f(u(t)),$
and the Banach embedding
$j:\mathscr{H}\rightarrow L^{2}(S^{1};V),\quad j(u)(t):=u(t).$
Since $\mathscr{L}:\mathscr{H}\rightarrow L^{2}(S^{1};V)$ is a linear isomorphism, $N:L^{2}(S^{1};V)\rightarrow L^{2}(S^{1};V)$ is continuous and $j:\mathscr{H}\rightarrow L^{2}(S^{1};V)$ is compact, the operator $\mathscr{F}:\mathscr{H}\rightarrow\mathscr{H}$ given by
$\mathscr{F}(u)=0,$ (6)
in the sense that a function $u\in\mathscr{H}$ is a solution to (4) if and only if it satisfies (6). In what follows, (4) will be called the operator equation associated with the system (4). Gracia-Azpetia et al. demonstrate the applicability of the $G$-equivariant Leray-Schauder degree to (6) by proving that the assumptions ($A_{1}$)($A_{4}$) imply that the nonlinear operator (5) is a completely continuous $G$-equivariant field, differentiable at the origin $0\in\mathscr{H}$ with
$\mathscr{A}u:=\operatorname{Id}-\mathscr{L}^{-1}(\beta^{2}Au-u),\quad\mathscr{A}:=D\mathscr{F}(0):\mathscr{H}\rightarrow\mathscr{H}.$
In order to effectively make use Lemma 1.1 to determine the existence of non-trivial solutions to (1), we must derive a practical formula for the computation of the local bifurcation invariant $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))\in A(G)$. Our first step in this direction will be to take advantage of the $G$-isotypic decomposition (3) to collect spectral data related to the $G$-equivariant linear operator $\mathscr{A}:\mathscr{H}\rightarrow\mathscr{H}$. By Schur’s Lemma, one has
$\mathscr{A}(\mathscr{H}_{m,j})\subset\mathscr{H}_{m,j},\quad m\in{\mathbb{N}}\cup\{0\},\;j\in\{0,1,\ldots,r\},$